TOPIC: CIRCLES DETERMINED BY THREE CONDITIONS
Write the equation (standard and general form)of the circle
1. through (-1, 2), (1, 1), and (3, 2)

Answer by richwmiller(17219)

Too many problems.
Limit of 5 problems per day
no similar problems
submit one problem at a time.
1) (x-1)^2+(y-7/2)^2 = 25/4
(x-a)^2+(y-b)^2 = c
(-1, 2), (1, 1), and (3, 2)
(-1-a)^2+(2-b)^2 = c
(1-a)^2+(1-b)^2 = c
(3-a)^2+(2-b)^2 = c
solve for a,b and c
a = 1, b = 7/2, c = 25/4